OpenStudy (anonymous):

Here is a distribution of 6 observations, sorted in ascending order: 4.7, 9.2, 10.9, 12.3, 16.6, 18.3 The mean of this distribution is 12. What is the value of ? 6 sigma (xi-x with the line above it)

OpenStudy (amistre64):

sigma as in: o~ ?

OpenStudy (anonymous):

no the one that looks like a capital E

OpenStudy (amistre64):

\[\sum(xi-\bar x)\]

OpenStudy (anonymous):

Yes!

OpenStudy (zarkon):

0

OpenStudy (amistre64):

i take it this is more properly: \[\sum_{i=1}^{6}(x_i-\bar x)\]

OpenStudy (anonymous):

Yes that is the correct equation

OpenStudy (amistre64):

it should, as zarkon states, be zero just subtract the mean from your vaoues and add up the totals

OpenStudy (amistre64):

values that is

OpenStudy (amistre64):

4.7 -12 = -7.3 9.2 -12 = -2.8 10.9 -12 = -1.1 12.3 -12 = .3 16.6 -12 = 4.6 18.3 -12 = 6.3 ----------------

OpenStudy (zarkon):

\[\sum_{i=1}^{n}(x_i-\bar x)\] \[=\sum_{i=1}^{n}x_i-\sum_{i=1}^{n}\bar x\] \[=\sum_{i=1}^{n}x_i-n\bar x\] \[=\sum_{i=1}^{n}x_i-n\frac{\sum_{i=1}^{n}x_i}{n}\] \[=\sum_{i=1}^{n}x_i-\sum_{i=1}^{n}x_i=0\]